OpenStudy (anonymous):

What is the LCM of 24a^3b and 36ab^2?

OpenStudy (mathstudent55):

You need to list the prime factors of each expression: \(24a^3b = 2^3 \times 3 \times a^3 \times b\) \(36ab^2 = 2^2 \times 3^2 \times a \times b^2 \)

OpenStudy (mathstudent55):

For the LCM, you need common factors and not common factors with highest exponent: \(LCM = 2^3 \times 3^2 \times a^3 \times b^2 \) \(LCM = 72a^3b^2\)